# MMEvol: Empowering Multimodal Large Language Models with Evol-Instruct

### v1

@@ Abstract
The development of Multimodal Large Language Models (MLLMs) has seen significant advancements. However, the quantity and quality of multimodal instruction data have emerged as significant bottlenecks in their progress. Manually creating multimodal instruction data is both time-consuming and inefficient, posing challenges in producing instructions of high complexity. Moreover, distilling instruction data from black-box commercial models (e.g., GPT-4o, GPT-4V) often results in simplistic instruction data, which constrains performance to that of these models. The challenge of curating diverse and complex instruction data remains substantial. We propose MMEvol, a novel multimodal instruction data evolution framework that combines fine-grained perception evolution, cognitive reasoning evolution, and interaction evolution. This iterative approach breaks through data quality bottlenecks to generate a complex and diverse image-text instruction dataset, thereby empowering MLLMs with enhanced capabilities. Beginning with an initial set of instructions, SEED-163K, we utilize MMEvol to systematically broaden the diversity of instruction types, integrate reasoning steps to enhance cognitive capabilities and extract detailed information from images to improve visual understanding and robustness. To comprehensively evaluate the effectiveness of our data, we train LLaVA-NeXT using the evolved data and conduct experiments across 13 vision-language tasks. Compared to the baseline trained with seed data, our approach achieves an average accuracy improvement of 3.1 points and reaches state-of-the-art (SOTA) performance on nine tasks. The project page is available at https://rainbowluocs.github.io/MMEvolWeb.
@@ 1 Introduction
Benefiting from the success of Large Language Models (LLMs)  (Touvron et al., 2023; Bai et al., 2023a; Zheng et al., 2024; Young et al., 2024; Tao et al., 2024), Multimodal Large Language Models (MLLMs) have experienced rapid development over the past two years, becoming the preferred approach for various vision-language tasks (Liu et al., 2024b; a; Li et al., 2023b; Dong et al., 2023). By aligning visual encoders with LLMs through large-scale coarse-grained image-text pre-training, followed by small-scale fine-grained supervised fine-tuning (SFT), MLLMs have demonstrated impressive capabilities in numerous vision-language tasks (Bai et al., 2023b; Luo et al., 2024; Lu et al., 2024). Existing open-source image-text instruction data primarily consists of two components. One of these is sourced from human-generated image-text instructions  (Chen et al., 2024a). However, due to the challenges humans may face in producing high-complexity instructions, these data often share only a limited number of common directives and execute similar tasks, with the difficulty distribution skewed towards simple or moderate levels. The other component originates from distilled prompts constructed using questions directed at commercial models, such as GPT-4o and GPT-4V (Liu et al., 2024b; c; Chen et al., 2023). Given those image-text instructions must adhere to visual content constraints, the distilled instruction data tends to be simple visual question-answering and scene description tasks. This severely restricts the diversity and complexity of existing image-text instruction data. The bottleneck in the quantity and quality of instruction data leads to models trained on such data frequently struggling to follow complex and varied user-specified instructions or objectives. This limitation reduces their practicality and applicability in real-world scenarios. Therefore, developing an automated method capable of producing open-domain instructions (particularly more challenging ones) at a relatively low cost has become crucial for advancing the instruction tuning of MLLMs.
@@
Inspired by the Evol-Instruct method (Xu et al., 2023; Luo et al., 2023b; a), we introduce MMEvol, which leverages multiple iterative evolutions of advanced MLLMs to automatically and massively generate a variety of open-domain instructions across different difficulty levels, thereby enhancing the performance of MLLMs. Given that vision-language instruction data is constrained by visual content, we made several adjustments to the evolutionary prompting process, specifically optimizing it for vision-language-related tasks. This effort culminated in three evolutionary directions: fine-grained perceptual evolution, cognitive reasoning evolution, and interactive evolution. The MMEvol mechanism is summarized in Fig. 1, where each evolutionary cycle comprises two main steps: instruction evolution and instruction elimination. Our method randomly selects from fine-grained perceptual evolution, cognitive reasoning evolution, or interactive evolution to upgrade simple instructions into more complex and diverse directives or to create new instructions, thereby increasing diversity and complexity. Fine-grained perceptual evolution aims to fully exploit the visual information within images, providing new directives that involve more detailed objectives. Cognitive reasoning evolution seeks to enhance the visual operational reasoning steps of the instructions, thereby increasing their complexity. Interactive evolution is designed to improve the diversity of instructions by offering a wider variety of directive forms. These three operations are achieved by providing specific prompts to the MLLM. Given that instructions generated through evolution may occasionally fail, we employed an instruction eliminator to filter out unsuccessful directives, termed elimination evolution. This evolutionary process is repeated multiple times to obtain an adequate dataset of instructions encompassing various complexities.
Specifically, we adapt LLaVA-Instruct (Touvron et al., 2023) and ShareGPT4V (Chen et al., 2023), supplemented by a small amount of additional scientific and chart data to ensure diversity, leading to 163K seed instruction samples. We perform three rounds of instruction evolution on this seed data using the OpenAI GPT-4o mini API, yielding 447K evolved samples. We fine-tune the open-source LLaVA-NeXT (Liu et al., 2024a) model with this evolved data, comparing it with other methods across multiple visual-language benchmarks. Our method achieved state-of-the-art (SOTA) performance, validating the effectiveness and efficiency of MMEvol. Additionally, we conduct thorough qualitative analysis and ablation experiments to demonstrate the contribution of each component of our method. We hope the release of evolved data and code will help the community better understand the role and significance of using high-quality image-text instruction data to train more powerful MLLMs.
Our main contributions can be summarized as follows:
@@ 2 Method
In this section, we first introduce the curation of seed instruction data and then elaborate on the methodological details of MMEvol. Due to the space limitation, we simplify the seed data curation process and prompt templates. More details are in the Appendix D.
@@ 2.2 Methodological Details
The evolution of image-text instruction data is constrained by visual information, requiring evolved instruction data to be relevant to image content to avoid hallucinations. This makes the diversity evolution of image-text instructions particularly challenging. Additionally, the complexity evolution process of image-text instruction data often results in shallow reasoning phenomena, with MLLMs struggling to provide complex answers. As shown in Fig. 1, to address these issues and improve the success rate of evolution, we include carefully designed domains such as visual objects, atomic capabilities, visual manipulations, and instruction formats to standardize each instruction data format. The visual object domain includes visual objects in the images involved in the instruction data, implicitly constraining the evolution data and reducing visual hallucinations. We also summarize nine types of atomic capabilities involved in image-text instruction data to populate the atomic capability domain, aiming to enhance data diversity. Specifically, this includes five vision-centric capabilities: localization, reference, computation, optical character recognition (OCR), and existence judgment, and four language-centric capabilities: relation description, scene understanding, behavior prediction, and world knowledge association. The visual manipulation domain includes visual manipulation chains (Qi et al., 2024) for problem-solving, where each step of the visual manipulation is based on vision-centric atomic capabilities, explicitly defining the visual reasoning process to mitigate shallow reasoning. The instruction format domain specifies the interaction types of the instruction data. These adaptations enhance the diversity and complexity of image-text instruction data and improve the success rate of evolution.
@@ 3.1 Benchmarks
To comprehensively evaluate the effectiveness of our evolutionary method, we select 13 benchmarks, with their sources and tested skills illustrated in Table 1. MIA (Qian et al., 2024) is an open-domain instruction-following benchmark that thoroughly tests the model’s instruction-following abilities using extensive instruction data. MM-Self-Instruct (Zhang et al., 2024) is a novel visual reasoning benchmark that focuses on the model’s visual perception capabilities and performs common visual reasoning tasks encountered in daily life.
@@
Visual Perception BLINK (Fu et al., 2024), RWQA MMSInst  (Zhang et al., 2024)
@@ 3.2 Implementation Details
Model. We follow the architecture from LLaVA-NeXT, where a multimodal large model consists of three key components: an LLM for next token prediction, a visual encoder for extracting visual features, and an image-text projector to align the visual and text modalities. We use Llama3-8B-Instruct (Touvron et al., 2023) for ablation experiments. For comparisons with other methods, we switch to our previous SOTA settings with Llama3-8B-Instruct and Qwen2-7B-Instruct (Bai et al., 2023a). We adapt CLIP-ViT-L (Radford et al., 2021) for the visual encoder and use simple linear layers to bridge the image and text modalities.
Training Strategies. We conduct MMEvol training following widely used two-stage settings. Vision-Language Pre-training and Visual Instruction-tuning. The language models and ViT are separately pre-trained, while the projector is randomly initialized. To initially align the feature space between the visual and text modalities, we utilize the aligned dataset. Finally, we perform instruction tuning of the pre-trained model on visual language instruction datasets. Our experiments are conducted with 8$\times$×A100 GPUs and a global batch size of 128. We employ AdamW optimizer Loshchilov (2017) with learning rates $5\times 10^{-5}$ and $2\times 10^{-5}$ for aforementioned two stages respectively. Each stage is trained with one epoch with a 3% warmup strategy. Please refer to the Appendix B for more details.
@@ 3.4 Ablation Study
We conduct ablation studies on seven vision-language benchmarks to explore the effects of instruction evolution and elimination. As shown in Table 2, using instruction evolution alone significantly enhances data diversity and complexity, leading to an average performance gain of 3.8 points across multiple vision-language benchmarks. However, the absence of instruction elimination introduces harmful data from failed evolutions, which inevitably reduces the model’s resistance to hallucinations by 1.2 points on POPE (Li et al., 2023c). When both instruction evolution and instruction elimination are employed, instruction elimination filters out harmful data from failed evolutions, further improving the quality and density of evolved data and enhancing the model’s performance by 0.9 points on average, particularly improving resistance to hallucinations by 1.7 points, which aligns with our qualitative analysis results in Section 3.3.
@@
Data MMStar MathVistaM POPE AI2D MMEC MMMUV RWQA AVG.
SEED Data 30K 36.5 25.3 84.8 53.9 31.5 32.3 43.5 44.0
+ Evolution w/o Elimination 38.9 (+3.4) 27.3 (+3.0) 83.6 (-1.2) 54.7 (+0.8) 40.1 (+8.6) 34.4 (+0.9) 54.4 (+10.9) 47.6 (+3.8)
+ Evolution w Elimination 40.3 (+3.8) 28.6 (+3.6) 86.5 (+1.7) 55.2 (+1.3) 39.9 (+8.4) 35.3 (+3.0) 55.3 (+11.8) 48.7 (+4.7)
@@ 4 Related Work
Multimodal Large Language Models (MLLMs). In recent years, the rapid development of MLLMs has been driven by the success of large language models and the availability of diverse image-text instruction data on the internet. Models like LLaVA (Liu et al., 2024b) and InstructBLIP (Dai et al., 2024) connect visual encoders and large language models using simple connectors, enabling robust cross-task generalization through joint training on multi-source instruction data. DreamLLM (Dong et al., 2023) extends the boundaries of MLLMs by employing diffusion models for image generation. DEEM (Luo et al., 2024) further refines the MLLM architecture and enhances robustness by substituting visual encoders with diffusion models for feature extraction. Despite these advancements, performance remains constrained by data quality, limiting further progress. ALLaVA (Chen et al., 2024a) initially uses manually crafted instruction data and rewrites answers. However, this approach is expensive and relatively simplistic, depending heavily on annotators’ experience. MM Instruct (Liu et al., 2024c), on the other hand, uses advanced MLLMs to generate instruction data automatically. Yet, the complexity and diversity of this data are limited by predefined formats, failing to leverage the visual information available fully.
Instruction Evolution. WizardLM (Xu et al., 2023) introduces instruction evolution into LLMs, resulting in significant performance improvements. This approach is further advanced by WizardCoder (Luo et al., 2023b) and WizardMath (Luo et al., 2023a), which use instruction evolution to enhance the coding and mathematical reasoning abilities of LLMs. VILA2 (Fang et al., 2024) tries to use instruction evolution to self-bootstrap but achieves limited success due to shallow reasoning observed in multimodal models, generating a vast amount of data that is overly simplistic and with limited diversity, making it difficult for other models to utilize effectively. In contrast, we address this challenge and propose MMEvol, which iteratively enhances the diversity and complexity of instruction data through evolution on limited datasets, efficiently extracting useful visual information and endowing multimodal models with more powerful capabilities.

### v2

@@ Abstract
The development of Multimodal Large Language Models (MLLMs) has seen significant advancements with increasing demands in various fields (e.g., multimodal agents, embodied intelligence). While model-driven approaches attempt to enhance MLLMs capabilities through diverse architectures, the gains have become increasingly marginal. Conversely, data-driven methods, which scale up image-text instruction data, are more effective but face limited data diversity and complexity challenges. The absence of high-quality data constitutes a significant development barrier for MLLMs. To address the data quality bottleneck, we propose MMEvol, a novel multimodal instruction data evolution framework. This framework iteratively improve data quality through a refined combination of fine-grained perception, cognitive reasoning, and interaction evolution, generating a more complex and diverse image-text instruction dataset that empowers MLLMs with enhanced capabilities. Beginning with an initial set of instructions, SEED-163K, we utilize MMEvol to systematically broaden the diversity of instruction types, extend visual reasoning steps to improve cognitive reasoning abilities, and thoroughly explore fine-grained information within images to enhance visual understanding and robustness. To comprehensively evaluate the effectiveness of our approach, we conduct extensive qualitative analysis and quantitative experiments across 13 vision-language tasks. Compared to baseline models trained with the initial seed data, the results demonstrate that our method achieves an average accuracy improvement of 3.1 percentage points. Furthermore, our approach reaches state-of-the-art (SOTA) performance in nine tasks using significantly less data compared to state-of-the-art models. The project page is available at https://rainbowluocs.github.io/MMEvolWeb.
@@ 1 Introduction
“The True Acquisition of Knowledge Lies in Grasping the Most Subtle Details.”
Aristotle, circa 4th century BCE
Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; a; Li et al., 2023b; Dong et al., 2023; Sun et al., 2023b; Dai et al., 2024; Luo et al., 2024; Qi et al., 2024) have seen rapid development over the past two years and have become the preferred approach for various vision-language tasks (Kembhavi et al., 2016; Fu et al., 2024; Zhang et al., 2024a; Qian et al., 2024). By aligning visual encoders (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023a) with LLMs (Touvron et al., 2023; Bai et al., 2023; Lu et al., 2024; Young et al., 2024; Tao et al., 2024), and employing large-scale coarse-grained image-text pre-training (Zhu et al., 2024; Schuhmann et al., 2022; 2021) followed by small-scale instruction-tuning (Chen et al., 2024a; Liu et al., 2024b), MLLMs have demonstrated impressive capabilities across numerous vision-language tasks and are widely applied in many domains (e.g., multimodal agents, embodied intelligence). Model-driven approaches (Luo et al., 2024; Liu et al., 2024a; Tong et al., 2024; Zhang et al., 2024b) aim to integrate knowledge from images and text more efficiently by designing different network architectures to improve MLLMs performance. However, their effectiveness is diminishing due to the lack of high-quality data, and redundant model designs struggle to realize their potential and push the boundaries of model intelligence. In contrast, data-driven methods (Liu et al., 2024b; Chen et al., 2024a; Yu et al., 2023; Liu et al., 2024c; Fang et al., 2024; Chen et al., 2023), while more effective, often produce data with limited diversity and complexity. The lack of high-quality data remains a significant obstacle in the development of MLLMs. Therefore, the need of developing automated methods that can generate more challenging and diverse instructional data at a relatively low cost is urgent for empowering MLLMs with enhanced capabilities.
@@
Analysis of existing data-driven methods for generating image-text instruction data reveals three common limitations: 1) Limited instruction diversity. Manually annotated instructions are constrained by the cognitive limitations of annotators, while model-generated instructions are limited by template presets, making it difficult to meet the diverse task requirements of the real world. This restricts the instruction-following ability of MLLMs. 2) Limited instruction complexity. Manual annotations often result in instructions of simple or moderate complexity, and automatically generated instructions tend to be brief and lacking in visual reasoning steps, which limits the model’s ability to handle complex tasks. 3) Insufficient alignment granularity. Both manually and model-generated instructions primarily focus on common objects, neglecting rare or small objects, resulting in limited granularity in image-text alignment. This affects the model’s visual perception robustness and resistance to hallucinations.
To address these limitations, we propose MMEvol, a novel method that utilizes advanced MLLMs for iterative evolution. This method automatically generates various types of open-domain instructions on a large scale, covering different difficulty levels to enhance the performance of MLLMs. Given that visual-language instruction data are constrained by visual content, the data generated through multiple iterations with Evol-Instruct (Xu et al., 2023; Luo et al., 2023a; b) tend to include simple restatements and data unrelated to visual content, making deep and broad evolution challenging. Therefore, we have made several adjustments to the evolution prompting process, ultimately developing an image-text instruction evolution paradigm. These adjustments include a more refined image-text instruction data paradigm and the definition of three evolution directions: fine-grained perception evolution, cognitive reasoning evolution, and interaction evolution. The MMEvol mechanism is summarized in Fig. 1, with each evolution cycle comprising two main steps: instruction evolution and instruction elimination. Instruction evolution randomly selects one of fine-grained perception evolution, cognitive reasoning evolution, or interaction evolution, upgrading simple instructions to more complex or diverse ones. Specifically, fine-grained perception evolution aims to leverage visual information in images to generate data with more detailed information; cognitive reasoning evolution prolongs the visual operation reasoning steps of instructions to increase their complexity; and interaction evolution aims to enhance instruction diversity by providing a wider variety of instruction forms. To account for occasional failures in evolved instructions, we use instruction elimination to filter out failed evolution. MMEvol repeats the instruction evolution and elimination processes multiple times to obtain a complex instruction dataset containing various instruction forms.
@@
To validate the effectiveness of MMEvol, we perform three rounds of evolutionary iterations on 163K seed data, leading to 447K evolved samples. We fine-tuned the open-source LLaVA-NeXT (Liu et al., 2024a) model with these evolved data and compared it with other advanced methods across 13 vision-language benchmarks. Our method achieves state-of-the-art (SOTA) performance, demonstrating the effectiveness and efficiency of MMEvol. Additionally, we conduct detailed qualitative analysis and ablation experiments to showcase the contribution of each component of our method. We hope that the released evolutionary data and code will assist the community in understanding that using a small amount of high-quality image-text instruction data is far more critical than training MLLMs with large-scale low-quality image-text instruction data.
Our main contributions can be summarized as follows:
@@ 2 Method
In this section, we first introduce the curation of seed instruction data and then elaborate on the methodological details of MMEvol. Due to the space limitation, we simplify the seed data curation process and prompt templates. More details can be found in the Appendix D.
@@ 2.2 Methodological Details
The evolution of image-text instruction data is constrained by visual information, requiring evolved instruction data to be relevant to image content to avoid hallucinations. This makes the diversity evolution of image-text instructions particularly challenging. Additionally, the complexity evolution process of image-text instruction data often results in shallow reasoning phenomena, with MLLMs struggling to provide complex answers. As shown in Fig. 1, to address these issues and improve the success rate of evolution, we include carefully designed domains such as visual objects, atomic capabilities, visual manipulations, and instruction formats to standardize each instruction data format. The visual object domain includes visual objects in the images involved in the instruction data, implicitly constraining the evolution data and reducing visual hallucinations. We also summarize nine types of atomic capabilities involved in image-text instruction data to populate the atomic capability domain, aiming to enhance data diversity. Specifically, this includes five vision-centric capabilities: localization, reference, computation, optical character recognition (OCR), and existence judgment, and four language-centric capabilities: relation description, scene understanding, behavior prediction, and world knowledge association. The visual manipulation domain includes visual manipulation chains for problem-solving, where each step of the visual manipulation is based on vision-centric atomic capabilities, explicitly defining the visual reasoning process to mitigate shallow reasoning. The instruction format domain specifies the interaction types of the instruction data. These adaptations enhance the diversity and complexity of image-text instruction data and improve the success rate of evolution.
@@ 3.1 Benchmarks
To comprehensively evaluate the effectiveness of our evolutionary method, we select 13 benchmarks, with their sources and tested skills illustrated in Table 1. MIA (Qian et al., 2024) is an open-domain instruction-following benchmark that thoroughly tests the model’s instruction-following abilities using extensive instruction data. MM-Self-Instruct (Zhang et al., 2024a) is a novel visual reasoning benchmark that focuses on the model’s visual perception capabilities and performs common visual reasoning tasks encountered in daily life.
@@
Visual Perception BLINK (Fu et al., 2024), RWQA MMSInst  (Zhang et al., 2024a)
@@ 3.2 Implementation Details
Model. We follow the architecture from LLaVA-NeXT, where a multimodal large model consists of three key components: an LLM for next token prediction, a visual encoder for extracting visual features, and an image-text projector to align the visual and text modalities. We use Llama3-8B-Instruct (Touvron et al., 2023) for ablation experiments. For comparisons with other methods, we switch to our previous SOTA settings with Llama3-8B-Instruct and Qwen2-7B-Instruct (Bai et al., 2023). We adapt CLIP-ViT-L (Radford et al., 2021) for the visual encoder and use simple linear layers to bridge the image and text modalities.
Training Strategies. We conduct MMEvol training following widely used two-stage settings. Vision-Language Pre-training and Visual Instruction-tuning. The language models and ViT are separately pre-trained, while the projector is randomly initialized. To initially align the feature space between the visual and text modalities, we utilize the aligned dataset. Finally, we perform instruction tuning of the pre-trained model on visual language instruction datasets. Our experiments are conducted with 8$\times$×A100 GPUs and a global batch size of 128. We employ AdamW optimizer (Loshchilov, 2017) with learning rates $5\times 10^{-5}$ and $2\times 10^{-5}$ for aforementioned two stages respectively. Each stage is trained with one epoch with a 3% warmup strategy. Please refer to the Appendix B for more details.
@@ 3.4 Ablation Study
We conduct ablation studies on seven vision-language benchmarks to explore the effects of instruction evolution and elimination. As shown in Table 2, different evolution process can be orthogonally superimposed on each other to continuously enhances data diversity and complexity. leading to an average performance gain of 3.8 points across multiple vision-language benchmarks. However, the absence of instruction elimination introduces harmful data from failed evolutions, which inevitably reduces the model’s resistance to hallucinations by 1.2 points on POPE  (Li et al., 2023c). When both instruction evolution and instruction elimination are employed, instruction elimination filters out harmful data from failed evolutions, further improving the quality and density of evolved data and enhancing the model’s performance by 0.9 points on average, particularly improving resistance to hallucinations by 1.7 points, which aligns with our qualitative analysis results in Section 3.3.
@@
FP-Evol I-Evol CR-Evol I-Elim MMStar MathVistaM POPE AI2D MMEC MMMUV RWQA AVG.
36.5 25.3 84.8 53.9 31.5 32.3 43.5 44.0
37.3 (+0.8) 25.6 (+0.3) 85.0 (+0.2) 54.2 (+0.3) 33.0 (+1.5) 32.5 (+0.2) 46.7 (+3.2) 44.9 (+0.9)
38.2 (+1.7) 26.2 (+0.9) 83.8 (-1.0) 54.5 (+0.6) 35.6 (+4.1) 32.9 (+0.6) 48.9 (+5.4) 45.8 (+1.8)
38.9 (+3.4) 27.3 (+3.0) 83.6 (-1.2) 54.7 (+0.8) 40.1 (+8.6) 34.4 (+0.9) 54.4 (+10.9) 47.6 (+3.8)
40.3 (+3.8) 28.6 (+3.6) 86.5 (+1.7) 55.2 (+1.3) 39.9 (+8.4) 35.3 (+3.0) 55.3 (+11.8) 48.7 (+4.7)
@@ 4 Related Work
Multimodal Large Language Models (MLLMs). MLLMs have rapidly advanced in recent years due to the success of Large Language Models (LLMs) and the availability of diverse image-text instruction data from the internet. LLaVA (Liu et al., 2024b) and MiniGPT-4 (Zhu et al., 2023) have demonstrated strong cross-task generalization by integrating visual encoders with large language models through simple connectors and training on instruction data. LLaVA-NeXT (Liu et al., 2024a) has significantly enhanced visual perception by employing dynamic resolution techniques. Cambrain-1 (Tong et al., 2024) has improved model robustness through visual encoder routing, though it incurs higher training costs. DEEM (Luo et al., 2024) simplifies model architecture and enhances robustness by using diffusion models to extract visual features instead of traditional visual encoders. Subsequent work (Wang et al., 2024b; Zhou et al., 2024; Xie et al., 2024) following DEEM combine diffusion models with LLMs to further enhance generative and understanding capabilities of MLLMs. However, these models still face challenges related to the quantity and quality of data, which limit performance improvements further.
Image-text Instruction Data Construction. LLaVA (Liu et al., 2024b) has improved model capabilities by utilizing LLaVA-Instruct (Liu et al., 2024b), a dataset labeled by advanced LLMs. However, this approach does not fully exploit visual information and have limited instruction types. ALLaVA (Chen et al., 2024a), by manually crafting and rewriting instruction data, offers greater variety but suffers from high manual labeling costs, inefficiency, and overly simplistic problems. MMInstruct (Liu et al., 2024c) generates instruction data automatically with advanced MLLMs, but the instruction complexity and diversity are constrained by predefined formats, failing to fully exploit effective visual information. VILA2 (Fang et al., 2024) has generated extensive data through instruction evolution but lacks complexity and variety, limiting its utility for other models. In contrast, we address this challenge and propose MMEvol, which iteratively enhances instruction diversity and complexity through instruction evolution on limited data, aiming to extract more usable visual information and endow MLLMs with more powerful capabilities.